Fermi National Laboratory
 

Accelerator Update for July 12, 2006 - July 14, 2006

 

Wednesday July 12

The day shift began with Operations monitoring store 4822 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with NuMI off due to a horn cooling problem, and with MiniBooNE taking beam.

Operations established beam to the Meson MTest experiment at 9 AM.

A Recycler expert began a RF study period at 9:33 AM.

At 10:16 AM, a Linac RF station (KRF4) tripped off. Operators contacted Linac technicians.

At 10:35 AM, KRF2 & 3 tripped off on a LCW flow indication.

At 10:39 AM, Operations reported that two Booster quadrupole power supplies (B:Q2 & B:Q5) were having problems. Booster experts investigated.

Linac techs had KRF3 on line by 10:42 AM, but KRF2 was still off. Operations contacted a water tech.

At 11:05 AM, Operations reported that Linac was back up and running, but that Booster was still down due to B:Q2 & 5.

At 11:33 AM, Operations reported that store 4822 aborted due to vacuum valves at TeV sector B4 closing. Vacuum experts investigated and found that an enable card had failed.

Pbar resumed stacking 1:05 PM.

Recycler studies finished at 1:11 PM.

Operations began preparing Recycler at 1:15 PM to stash antiprotons.

Booster experts reported at 1:24 PM that after they replaced a power supply for B:Q2, B:Q5 began operating normally.

At 2 PM, an Accumulator RF station (ARF4) tripped off. Experts investigated.

The stash ended and Pbar resumed stacking at 2:26 PM.

At 2:45 PM, Operation reported that B:Q2 & 5 were again acting oddly. Experts investigated.

A Recycler expert began a MI-60 study at 3:19 PM.

Booster resumed sending beam at 3:21 PM, but then went down only eight minutes later. At 3:51 PM beam returned. Experts reported that they found and repaired a problem with a circuit breaker for the B2 & 5 power supplies.

Vacuum experts reported at 4:17 PM that all the valves in TeV sector B4 were open and the vacuum looked good.

TeV experts began a study period at 6:50 PM.

MTest finished using beam around 8:30 PM.

The Recycler began preparing to stash at 8:44 PM.

At 8:57 PM, the P1 transfer line between the TeV and MI suffered a vacuum burst. Operators closed vacuum valves to halt the cascade.

Operations put the TeV into shot setup at 10:30 PM.

Thursday July 13

The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operation established store 4825 at 12:47 AM with an initial luminosity of 65.7E30.

The Recycler began preparing to stash at 2:33 AM. The stash ended and Pbar resumed stacking at 3:10 AM.

At 6:21 AM, the H- Source tripped off due to vacuum valves closing.

The Recycler began preparing to stash at 7:23 AM. The stash ended and Pbar resumed stacking at 7:49 AM.

Operations established beam to MTest at 9:01 AM.

MI experts began aperture scans at 9:50 AM.

At 11:25 AM, LRF5 tripped off. Operators investigated and then called experts.

At 11:44 AM, store 4825 aborted due to the TeV sector B4 vacuum valves. Operators contacted vacuum experts.

The Linac experts adjusted the PA tuners and had the LRF5 back up and beam running by 12 AM.

The Recycler began preparing to stash at 12:30 PM.

Operations began preparing the TeV for an access at 12:35 PM.

Operations prepared CDF and D0 for access at 1:01 PM.

The stash ended at 1:07 PM.

Operations prepared Linac for access at 1:08 PM.

Operations began preparing the Pbar AP1 transfer line for access at 1:32 PM.

The AP1 access ended at 2:37 PM.

The TeV came out of access at 2:53 PM.

Operations reported at 3:11 PM that experts changed out a CIA crate at B4.

Linac came out of access and beam resumed at 3:13 PM.

Operations reported the following status at 3:25 PM:

  • The TeV turn on is in progress
  • The TeV vacuum work continues
  • The #2 Tevatron Electron Lens (TEL) is conditioning
  • TeV RF work continues
  • Linac is running beam
  • A Booster expert had placed the Booster in dump mode so he could work on the chopper
  • Booster RF work continues
  • MI is ramping
  • MiniBooNE enclosures MI-2 A&B are in access
  • Pbar is ready and waiting for beam

Operations established beam to the MI dump at 3:40 PM

Operations established beam to the Accumulator at 3:44 PM.

MiniBooNE came out of access at 3:46 PM.

Operations established beam to MiniBooNE at 4:23 PM.

At 5:49 PM, Operations tried to reset vacuum cards at TeV sector E3, B4, and A0, but the A0 reset didn’t take. Operators called in experts.

Operations put the TeV through a dry squeeze at 6:40 PM.

At 7:01 PM, BRF8 tripped off for the first of six times in two and a half hours. An expert investigated.

The Recycler began preparing to stash at 8:18 PM. MTest decided to quit for the night.

At 8:41 PM, B:Q2 & 5 tripped off again. Operations called in experts.

The stash ended and Pbar resumed stacking at 9:03 PM.

Booster experts came in and managed to get the supplies back on by 9:57 PM. Pbar resumed stacking and MiniBooNE resumed taking beam.

At 10:42 PM, Operations turned off the Booster beam for five minutes to troubleshoot B:Q2 & 5.

At 11:09 PM, B:Q2 & 5 tripped off again. A Booster expert stated that he thought the problem was with a wall breaker, but that repair would take a couple of hours. He made a phase adjustment to try and get through the night.

Friday July 14

The midnight shift began with Operations monitoring the stash, with vacuum techs troubleshooting the B4 vacuum controls and readbacks, and with Booster holding off beam to Pbar and MiniBooNE.

Pbar and MiniBooNE resumed taking beam at 12:05 AM.

Operations reported at 12:49 AM that they had readbacks for TeV sector E4, but not for A0.

Operations reported at 1:33 AM that all the TeV vacuum valves and controllers were working.

At 2:05 AM, Operations reported that pulses of beam to Booster were occasionally missing. This was causing beam position problems in MiniBooNE.

Operations put the TeV into a wet squeeze at 2:08 AM.

Operations put the TeV into shot setup around 2:43 AM.

At 4:25 AM, a TeV separator sparked (C49V).

Operations established store 4828 at 4:39 AM with an initial luminosity of 91.73E30. The TeV separator spark hurt the luminosity.

The Recycler began preparing to stash at 5:10 AM.

A Booster expert began working on the 400 MeV chopper around 5:11 AM.

At 5:16 AM, the Booster West Anode Power Supply (WAPS) tripped due to a BRF15 LCW flow problem. Operators had them both back on within twenty minutes.

The stash ended and Pbar resumed stacking at 6:11 AM.

At 6:53 AM, Operations reported that the LCW temperatures at TeV sectors A1 & 2 were high. An expert investigated and discovered that the pond pump was off. He managed to get the pump back on, but later reported that the heat exchanger for A1 was plugged.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

Machine Reports

Linac

Last Week

  • Running on the H- Ion Source
  • Beam Current 35 ma
  • Beam position problems have varied throughout the week being low to moderate

Downtime 2.39 Hours With 25 Occurrences

  • 98.6% Availability

Major Sources of Downtime

  • .78 hr. KRF4 Filament timeout
  • 16 min. Radiation monitor read-backs failed
  • LRF5 had sparking problem across PA Anode

Problems

  • Source decayed by 4 mA over week end and had to be tuned back
  • We made an access to repair NMR probe in spectrometer magnet and the repair failed

Booster

Last Week

  • Stacking: 8 -10 turns with ~ 3.5 to 4E12 per batch
  • MiniBooNE; 10 – 11 TURNS
  • NUMI: off

Studies

  • Tuning: Injection/capture – RF timing and paraphase
  • Transition crossing – modified phase jump control
  • Gamma-t
  • Preaccelerator notching – continues on $1D cycles

Downtime: 8 hours - many lost pulses

  • 400 MeV chopper – a problem with the off tube misfiring
  • B:Q2 and Q5 (400 MeV power supply) regulation issues (looks like a bad phase)

Next Week

  • Continue to work on pre-acc notching $1D
  • Push higher intensity – injection issues
  • Damper work – modes 47 & 2 still too high
  • Continue Gamma –t
  • Reposition collimators – orbit changes

Main Injector

MI operation

  • Shot to the TeV
    • Pbar coalescing eff. 86 - 88%
  • Shot to the RR
    • Beam loss @ 3 sec was gone
      • H Tune was increased by 0.003
  • Stacking
    • Intensity @injection ~ 7.5E12
    • Intensity @ target ~ 6.5E12 (Acc. eff. ~ 86%)

BPM upgrade

  • MI20, 30 & 40 crate repair - complete
  • MI60 South electronics installation - complete
  • Orbit closure study for reverse proton – continue
  • MI60 North electronics installation - next week?

MI studies

  • Aperture scan – continue
  • Slip stacking with one frequency off - next week

Antiproton Source

Stacking Observations

  • Present average stacking rate is 14.2 mA/hr. This is 6% better than last week and about 20% below the best week of stacking
  • Beam on target averaged about 6.0E12 POT this week. This is about 25% below the best week of stacking
  • Overall production efficiency averaged 14.2 antiprotons per 106 POT. This is 12% below the production efficiency during the best week of stacking
  • Debuncher 1st turn yield increased by 5-10% after repairing D:QS732. It is now higher than that of the best week of stacking. The antiproton yield into the Debuncher increased by approximately the same amount

Problems, Issues, and Solutions

  • Bad news
    • Core transverse cooling PLC failure -- Fixed
    • A:LQ contactor failure × 2 -- Fixed
    • 120 GeV beamline BPM problems => Overthruster OFF
    • Overthruster off for AP2 – Not Fixed
    • Debuncher kicker tank 201 still OOC
  • Good News
    • D:QS732 repaired => Debuncher injected beam up by 5-10%, AP50 beam position improved
    • Core VB1 fixed and back in service.
    • 10 dB gain added to core 4-8 GHz ∆p cooling

Recycler

Bunch Unevenness in Extraction with Momentum Mining

  • Caused by potential well distortion
  • Causes large variation in bunch intensity and numbers of antiprotons per transfer
  • Proof based on cavity fan back voltage
  • Made more apparent by small momentum spread
  • Can cause stability problems
  • Solution
    • Adaptive feed-forward correction
      • Demonstrated successfully on July 12
      • Fast response compared to speed of RF manipulation
      • Takes care of beam loading automatically

Status of Bucket Barrier Studies and Auto Correction

  • Success!
  • Works with moving buckets
  • Need to integrate into LLRF system
  • Need to measure noise and effect on longitudinal emittance
  • Need to test on complex waveforms
  • User interface and maintenance
  • In operation within 1.5 months?

Tevatron

Store Summary

Store

Avg Init. Lum (E30)

Avg Deliv’d

Lum

(nb -1)

Time (hr)

How did it end?

Comments

4825

65.4

1471

10.8

Abort

B4 Vacuum valve problems;

4822

105.6

965

3.2

Abort

Low beta tune small changes; BV8U vacuum valve control problem;

4821

157.5

2899

7.8

Abort

Tune changes for ramp and squeeze; SVX abort;

4818

151.1

6593

36.2

Intentional

TRF8 trip;

4816

139.6

4216

24.0

Intentional

Tune changes for ramp and squeeze; TRF1 trip; E1 cryo vacuum problem

Machine Maintenance and Studies

  • Machine tuning ups and optimization
    • Tune optimizations for ramp and squeeze
  • Reverse injection closure test for new MI BPM system
    • T117 needed modification
  • TEL2 tuning and conditioning, modulator connections
  • Orbit stabilization program upgrade

Problems

  • RF1 tripped and had an adverse impact upon luminosity, abort gap losses & beam lifetime.
  • C4 air compressor fail
  • E1 cryo system vacuum problem
  • SVX abort;
  • B4 Vacuum valve problem
  • D49V target motor drive card fail
  • Vertical flying wire problem

Study Requests

  • Proton bumper timing adjustment
    • Between stores (1hr)
  • A0 beam abort position adjustment
    • EOS (1hr)
  • Tune adjustment for squeeze
    • Between stores (2hrs)
  • TEL2 orbit and timing tuning, modulator test
    • Available proton only times
  • Digital Tune Monitor with VTICK (EOS 30min)
  • Helix orbit tuning

Switchyard/MTest

Last Week

  • T927 (CMS Pixel) started in proton mode 9am-9pm
  • Slow spill length modified from 1 sec to 4 sec at T927’s request previous week, request for ~100K in 120 GeV proton mode
  • ~80 hours requested, ~46 received ~60%
  • T927 scheduled to continue until July 24

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/14/2006    email Fermilab